# Explanation Evaluation

### v1

@@ Appendix E Crowdsourced Explanations for BCOPA
Explanation: Taking sleeping pills is a direct cause of awakening, whereas a husband snoring is an indirect cause.
Overall: 1.6 Well-written: Yes
Supports: a) Related: Yes
Factual: No
New info: 1.4
Unn. info: No
Contrastive: Yes
Crowdsourced (best) The man was bitten by mosquitoes. What was the cause of this? a) He fell asleep on his couch.
b) He went camping in the woods.
@@
Explanation: Sleeping on a couch would mean you are indoors where there are rarely mosquitoes. Mosquitoes are prevalent in wooded areas, so the man would be more likely to be camping in the woods if he was bit by mosquitoes.
Overall: 5.0 Well-written: Yes
Supports: b) Related: Yes
Factual: Yes
New info: 3.0
Unn. info: No
Contrastive: Yes
Crowdsourced (worst) I refilled my water bottle. What was the cause of this? a) I drank all the water in it.
b) I kept it in the fridge.
@@
Explanation: If you kept your water in the fridge, it would not need to be refilled.
Overall: 2.2 Well-written: Yes
Supports: a) Related: Yes
Factual: Yes
New info: 1.0
Unn. info: No
Contrastive: Yes
Table 5: Top-rated and bottom-rated examples of explanations for BCOPA in our test set. The ratings on the right are aggregated from five crowdworkers’ annotations. Well-written ranges from $1.0$ to $5.0$, and New info. from $0.0$ to $3.0$.
Figure 6: Explanation writing AMT form.
Figure 7: Explanation rating AMT form.

### v2

@@ Appendix D Crowdsourced Explanations for BCOPA
Generated (worst) The woman awakened. Cause? a) She took sleeping pills.     b) Her husband was snoring. Explanation: Taking sleeping pills is a direct cause of awakening, whereas a husband snoring is an indirect cause. Over.: 1 Well-wr.: yes
Supp.: a) Rel.: yes
Fact.: no
New i.: +
Unn. i.: no
Contr.: yes
@@
Crowdsourced (best) The man was bitten by mosquitoes. Cause? a) He fell asleep on his couch.     b) He went camping in the woods. Explanation: Sleeping on a couch would mean you are indoors where there are rarely mosquitoes. Mosquitoes are prevalent in wooded areas, so the man would be more likely to be camping in the woods, if he was bit by mosquitoes. Over.: 5 Well-wr.: yes
Supp.: b) Rel.: yes
Fact.: yes
New i.: $+$
Unn. i.: no
Contr.: yes
@@
Crowdsourced (worst) I refilled my water bottle. Cause? a) I drank all the water in it.     b) I kept it in the fridge. Explanation: If you kept your water in the fridge, it would not need to be refilled. Over.: 2 Well-wr.: yes
Supp.: a) Rel.: yes
Fact.: yes
New i.: +
Unn. i.: no
Contr.: yes
Table 6: Top- and bottom-rated example explanations for BCOPA in [Uncaptioned image]Acorn. New info. labels are denoted as follows: $+$ (ample), ++ (sufficient), + (some), - (none).
Model version Supp. Ovr. W-wr. Rel. Fact. New i. Un. i. Cntr. Avg.
gpt-3.5-turbo-0613 0.861 0.561 0.327 0.644 0.413 0.458 0.386 0.756 0.551
gpt-4-0613 0.873 0.624 0.346 0.673 0.427 0.476 0.397 0.812 0.578
gpt-4o-2024-05-13 0.859 0.624 0.409 0.666 0.441 0.485 0.416 0.794 0.587
gpt-4o-mini-2024-07-18 0.852 0.619 0.411 0.648 0.423 0.474 0.393 0.791 0.576
text-bison-001 0.846 0.588 0.344 0.629 0.407 0.468 0.411 0.649 0.543
gemini-1.0-pro 0.834 0.597 0.359 0.619 0.404 0.474 0.408 0.714 0.551
gemma-2-9b-it 0.836 0.553 0.360 0.634 0.403 0.407 0.368 0.748 0.539
gemma-2-27b-it 0.823 0.582 0.387 0.633 0.414 0.411 0.336 0.788 0.547
Meta-Llama-3.1-8B-*-# 0.802 0.593 0.396 0.630 0.366 0.439 0.415 0.717 0.545
Meta-Llama-3.1-70B-*-# 0.860 0.623 0.412 0.647 0.416 0.480 0.378 0.792 0.576
Meta-Llama-3.1-405B-*-# 0.860 0.637 0.405 0.644 0.424 0.482 0.385 0.804 0.580
Mixtral-8x7B-*-v0.1 0.834 0.586 0.356 0.622 0.397 0.455 0.358 0.703 0.539
Mixtral-8x22B-*-v0.1 0.848 0.629 0.386 0.645 0.428 0.472 0.428 0.808 0.580
Human 0.866 0.613 0.365 0.655 0.399 0.442 0.433 0.784 0.570
Table 7: Full results of the experiments comparing the difference in inter-annotator agreement between humans and with a random rater replaced by an LLM (§3.2). All values represent Krippendorff’s $\alpha$ averaged over 20 iterations. Extraction failures are excluded from analysis. Replace * with "Instruct" and # with "Turbo" in the model names.
Model version Supp. Ovr. W-wr. Rel. Fact. New i. Un. i. Cntr. Avg.
gpt-3.5-turbo-0613 0.846 0.627 0.409 0.696 0.621 0.611 0.383 0.786 0.622
gpt-4-0613 0.900 0.740 0.515 0.778 0.675 0.691 0.528 0.946 0.722
gpt-4o-2024-05-13 0.844 0.758 0.611 0.750 0.688 0.714 0.457 0.896 0.715
gpt-4o-mini-2024-07-18 0.853 0.708 0.620 0.700 0.628 0.678 0.463 0.888 0.692
text-bison-001 0.773 0.685 0.456 0.654 0.653 0.640 0.515 0.607 0.623
gemini-1.0-pro 0.787 0.681 0.497 0.613 0.549 0.671 0.451 0.648 0.612
gemma-2-9b-it 0.798 0.645 0.492 0.674 0.584 0.616 0.388 0.742 0.617
gemma-2-27b-it 0.769 0.626 0.574 0.668 0.603 0.537 0.191 0.867 0.604
Meta-Llama-3.1-8B-*-# 0.632 0.638 0.562 0.639 0.535 0.555 0.469 0.662 0.586
Meta-Llama-3.1-70B-*-# 0.847 0.732 0.632 0.703 0.652 0.680 0.311 0.886 0.680
Meta-Llama-3.1-405B-*-# 0.833 0.745 0.618 0.699 0.631 0.686 0.351 0.916 0.685
Mixtral-8x7B-*-v0.1 0.747 0.624 0.487 0.616 0.543 0.604 0.314 0.640 0.572
Mixtral-8x22B-*-v0.1 0.799 0.733 0.588 0.688 0.642 0.681 0.540 0.928 0.700
Table 8: Full results of the experiments measuring Spearman’s rank correlation between majority-voted human labels and LLM-generated ones (§3.3). Extraction failures are excluded from analysis. Replace * with "Instruct" and # with "Turbo" in the model names.